OpenStudy (anonymous):

in a group of 100 persons 60 like tea 7 like coffee. if x persons like both tea and coffee show 35 <= x <=6 0

OpenStudy (anonymous):

75 like coffee

OpenStudy (badhi):

Its obvious that \(x \leq 60\) since the group which like both tea and cofee is a subset of the group that like tea, and the group that like coffee. Thus, \(x \leq Min\{\text{no. of coffe people}, \text{no. of tea people} \}\) now from the venn diagram, \(60 + 75 -x \leq 100\) 

OpenStudy (triciaal):